Rohan have some books and some tables.If he puts 9 books into each table he will have two tables left vacant.If he puts 6 books on each table he will have 3 books left over.Find the numbers of books?
Answer by Edwin McCravy(20055)

Suppose there are B books and T tables

If he puts 9 books on each table he will have two tables left vacant.
B = 9(T-2)  

If he put 6 books on each table he will have 3 books left over .Find the numbers of books
6T+3 = B

6T+3 = 9(T-2)
6T+3 = 9T-18
  21 = 3T
   7 = T

6T+3 = 6(7)+3 = 42+3 = 45 = B

So there are 7 tables and 45 books.

Checking:  Putting 9 books on 5 tables takes all 45 books,
and there will be 7-5 = 2 tables with no books. 

Putting 6 books on all 7 tables takes 42 books,
and there will be 45-42 = 3 books left over.

Edwin